# Multivariate generalized Pareto distributions along extreme directions

### v3

@@ 3 Extreme directions
Upper tail dependence coefficients (Coles et al., 1999; Schlather and Tawn, 2002)
@@ 3.1 Definition and characterizations
Recall the exponent measure $\mu$ of $G^{*}$ as defined in (2.4) and $\mathbb{E}=[\boldsymbol{0},\boldsymbol{\infty})\setminus\left\{{\boldsymbol{0}}\right\}$. In this section, let $J\subseteq D$ be a fixed non-empty set and $\mathbb{E}_{J}:=\left\{{\boldsymbol{x}\in\mathbb{E}:\ x_{j}>0\text{ iff }j\in J}\right\}$. The collection $\left\{{\mathbb{E}_{J}:\;\emptyset\neq J\subseteq D}\right\}$ forms a partition of $\mathbb{E}$. The following definition was introduced in Simpson et al. (2020) in order to identify extreme directions of the random vector $\boldsymbol{X}$ defined as in (2.3).
@@
$\mathbb{R}_{n,J}^{\epsilon}$ =$=\left\{\boldsymbol{x}\in\mathbb{E}:\ \|\boldsymbol{x}\|>n,\;\boldsymbol{x}_{J}/n\in(\epsilon,\infty)^{\lvert J\rvert},\;\boldsymbol{x}_{D\setminus J}/n\in[0,\epsilon]^{\lvert D\setminus J\rvert}\right\},$ (3.2)
$\mathbb{C}_{n,J}^{\epsilon}$ =$=\left\{\boldsymbol{x}\in\mathbb{E}:\;\|\boldsymbol{x}\|>n,\;\boldsymbol{x}_{J}/\|\boldsymbol{x}\|\in(\epsilon,\infty)^{\lvert J\rvert},\;\boldsymbol{x}_{D\setminus J}/\|\boldsymbol{x}\|\in[0,\epsilon]^{\lvert D\setminus J\rvert}\right\},$ (3.3)
@@
The latter sets were introduced in Goix et al. (2017, 2016) for the $L_{\infty}$-norm. For $\epsilon>0$, the Borel set $\mathbb{R}_{1,J}^{\epsilon}$ is $\mu$-continuous as defined in Section 2. Hence we can make the following link with Proposition 3.2.
@@
$\boldsymbol{M}=\left(\max_{k\in R}\{a_{1k}Z_{k}\},\ldots,\max_{k\in R}\{a_{dk}Z_{k}\}\right).$
@@
$\chi_{J}=\mu\left(\left\{\boldsymbol{x}\in\mathbb{E}:\;\boldsymbol{x}_{J}\in(1,\infty)^{\lvert J\rvert}\right\}\right)=\sum_{\begin{subarray}{c}\bar{J}\subseteq D\\ \bar{J}\supseteq J\end{subarray}}\mu\left(\left\{\boldsymbol{x}\in\mathbb{E}:\;\boldsymbol{x}_{\bar{J}}\in(1,\infty)^{\lvert\bar{J}\rvert},\;\boldsymbol{x}_{D\setminus\bar{J}}\in[0,1]^{\lvert D\setminus\bar{J}\rvert}\right\}\right).$
We have $\chi_{J}>0$ if and only if there exists $\bar{J}\supseteq J$ such that the corresponding term on the right-hand side is positive, and by homogeneity of $\mu$, this is equivalent to $\bar{J}$ being an extreme direction. We find that $\chi_{J}>0$ if and only if $J$ is a subset of an extreme direction.
@@ 3.2 Density of an mgp distribution with multiple extreme directions
We provide density expressions for potential use in statistical inference on mgp models. As in Section 2.2, let $\boldsymbol{Y}$ be an mgp random vector with distribution $H=\operatorname{GP}(G)$ where $G$ is an mev distribution with Gumbel margins. Further, let $\boldsymbol{T}$ and $\boldsymbol{U}$ be its generators as defined in Section 2. With respect to a dominating measure $\upsilon$, defined as a sum of Lebesgue measures $\lambda_{m}$ of various dimensions $m\in D$, we express the density of $\boldsymbol{Y}$ in terms of the densities of $\boldsymbol{U}$ and $\boldsymbol{T}$. This will be helpful since the latter densities are known for several parametric families, for instance, the logistic and the Hüsler–Reiss model (Kiriliouk et al., 2018, Sections 7.1 and 7.2). Furthermore, in view of (2.9), the density of the exponent measure $\mu$ on $\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\boldsymbol{x}\nleqslant\boldsymbol{1}}\right\}$ is proportional to the mgp density, after appropriate marginal scaling; by homogeneity, this determines the exponent measure density everywhere.
@@
The proofs of Theorems 3.9 and 3.10 are identical to the ones of Theorems 12 and 13 in Rootzén et al. (2018). In the next section, we will propose a model for the $\boldsymbol{U}$-generator that leads to mgp distributions with general extreme directions and whose density can be computed by means of Theorem 3.9.
@@ 4.1 Definition and properties
$\ell_{J_{k}}^{(k)}(\boldsymbol{x}_{J_{k}})=\ell^{(k)}\left({\tilde{\boldsymbol{x}}}\right),\qquad\boldsymbol{x}_{J_{k}}\in[0,\infty)^{\lvert J_{k}\rvert},$ (4.3)
with $\tilde{\boldsymbol{x}}=\sum_{j\in J_{k}}x_{j}\boldsymbol{e}_{j}$ and $\boldsymbol{e}_{j}$ the $j$-th canonical unit vector in $\mathbb{R}^{d}$. If $J_{k}$ is a singleton, then $\ell_{J_{k}}^{(k)}$ is the identity function on $[0,\infty)$. The following theorem was proved in Stephenson (2003, Theorem 1) for the particular case where $J_{1},\ldots,J_{r}$ are all different.
@@
The random vector $\boldsymbol{M}$ in (4.1) follows an mev distribution $G_{\boldsymbol{M}}$ with unit-Fréchet margins and stdf
@@
In the context of anomaly detection, Chiapino et al. (2020) define a mixture model for the angular measure of $G_{\boldsymbol{M}}$ in terms of probability measures on the faces of the unit simplex. In contrast, the mixture model in Definition 4.1 is obtained by an aggregation of lower-dimensional mev distributions.
@@
as defined in Section 2.2. We will refer to $H$ as the mgp distribution associated with the mixture model $\boldsymbol{M}$ and to $\boldsymbol{Y}\sim H$ as the mgp random vector associated with the mixture model $\boldsymbol{M}$. For a non-empty set $J\subseteq D$, recall the set $\mathbb{A}_{J}$ and the projection $\pi_{J}$ defined in Eq. (2.1). For $k\in R$, since the only extreme direction of $\boldsymbol{Z}_{J_{k}}^{(k)}$ is $J_{k}$, Theorem 3.6 yields the existence of a random vector $\boldsymbol{U}^{(k)}$ taking values in $\mathbb{R}^{\lvert J_{k}\rvert}$ which satisfies (2.15) and
$\operatorname{\mathsf{E}}\left[{\max_{j\in J_{k}}\left\{{x_{j}\mathrm{e}^{U_{j}^{(k)}}}\right\}}\right]=\ell_{J_{k}}^{(k)}\left({\boldsymbol{x}_{J_{k}}}\right),\qquad\boldsymbol{x}_{J_{k}}\in[0,\infty)^{\lvert J_{k}\rvert}.$ (4.6)
@@ 4.2.1 Mixture logistic model
$\ell_{J_{k}}^{(k)}(y_{1},\ldots,y_{\lvert J_{k}\rvert})=\left({y_{1}^{1/\alpha_{k}}+\cdots+y_{\lvert J_{k}\rvert}^{1/\alpha_{k}}}\right)^{\alpha_{k}},$ (4.9)
for $\boldsymbol{y}\in[0,\infty)^{|J_{k}|}$, where $10<\alpha_{k}<$ 1. The only extreme direction of $\boldsymbol{Z}_{J_{k}}^{(k)}$ is $J_{k}$, so that, following Definition 4.1 and Remark 4.2, the mixture logistic model $\boldsymbol{M}$ is well-defined. Let $G_{\boldsymbol{M}}$ denote the cdf of $\boldsymbol{M}$; note that $G_{\boldsymbol{M}}$ is an mev distribution with unit-Fréchet margins. We compute its stdf using Theorem 4.4.
@@
In case the sets $J_{1},\ldots,J_{r}$ are all different, the stdf $\ell$ in (4.10) coincides with the one of the asymmetric logistic model introduced in Tawn (1990).
Next we construct a $\boldsymbol{U}$-generator of the mgp distribution $H=\operatorname{GP}(\tilde{G}_{\boldsymbol{M}})$, where $\tilde{G}_{\boldsymbol{M}}$ is the cdf of $\left({\ln M_{1},\ldots,\ln M_{d}}\right)$, an mev distribution with Gumbel margins. Let $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ be random vectors in $\mathbb{R}^{\lvert J_{1}\rvert},\ldots,\mathbb{R}^{\lvert J_{r}\rvert}$, respectively, such that
@@ 4.2.2 Mixture Hüsler–Reiss model
For $k\in R$, recall $J_{k}$ in (4.2) and suppose that $\boldsymbol{Z}_{J_{k}}^{(k)}$ has Hüsler–Reiss mev distribution (Hüsler and Reiss, 1989) with variogram matrix [^1] A matrix $\Gamma\in\mathbb{R}^{d\times d}$ is a variogram matrix if it is symmetric, has zero diagonal, and satisfies $\boldsymbol{v}^{\top}\Gamma\boldsymbol{v}<0$ for all $\boldsymbol{0}\neq\boldsymbol{v}^{\top}\perp\boldsymbol{1}$. To each positive definite covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$ is associated a variogram matrix $\Gamma\in\mathbb{R}^{d\times d}$ via $\Gamma_{st}=\Sigma_{ss}+\Sigma_{tt}-2\Sigma_{st}$ for $s,t\in D$. $\Gamma^{(k)}\in\mathbb{R}^{\lvert J_{k}\rvert\times\lvert J_{k}\rvert}$; see, e.g., Huser and Davison (2013). If $J_{k}$ is a singleton, the stdf $\ell_{J_{k}}^{(k)}$ of $\boldsymbol{Z}_{J_{k}}^{(k)}$ reduces to the identity function on $[0,\infty)$. Otherwise,
$\ell_{J_{k}}^{(k)}(\boldsymbol{y})=\sum_{j=1}^{\lvert J_{k}\rvert}y_{j}\,\Phi_{\lvert J_{k}\rvert-1}\left({\eta^{j,(k)}\left({\boldsymbol{y}}\right);\Sigma^{j,(k)}}\right),\qquad\boldsymbol{y}\in[0,\infty)^{\lvert J_{k}\rvert},$
where for $j\in J_{k}$, $\eta^{j,(k)}\left({\boldsymbol{y}}\right)$ is equal to $\left\{{\log(y_{j}/y_{s})+\Gamma^{(k)}_{js}/2}\right\}_{s=1,\ldots,\lvert J_{k}\rvert;s\neq j}$ with $y_{j}/0=\infty$ if $y_{j}>0$ and $\infty$ if $y_{j}=0$, where the matrix $\Sigma^{j,(k)}$ is defined by
$\Sigma^{j,(k)}:=\frac{1}{2}\left\{{\Gamma_{js}^{(k)}+\Gamma_{jt}^{(k)}-\Gamma_{st}^{(k)}}\right\}_{\begin{subarray}{c}s,t=1,\ldots,\lvert J_{k}\rvert\\ s,t\neq j\end{subarray}}\in\mathbb{R}^{\left({\lvert J_{k}\rvert-1}\right)\times\left({\lvert J_{k}\rvert-1}\right)}$
@@
To construct a generator of the mgp distribution $H=\operatorname{GP}(\tilde{G}_{\boldsymbol{M}})$ of the mixture Hüsler–Reiss model, let $k\in R$ and let $\tilde{\Sigma}^{(k)}\in\mathbb{R}^{\lvert J_{k}\rvert\times\lvert J_{k}\rvert}$ be a positive definite covariance matrix with variogram matrix $\Gamma^{(k)}$; see, e.g., Segers (2020) for a possible construction. Let $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ be random vectors in $\mathbb{R}^{\lvert J_{1}\rvert},\ldots,\mathbb{R}^{\lvert J_{r}\rvert}$, respectively, such that
@@
for $\boldsymbol{y}\in[-\infty,\infty)^{d}\setminus[-\infty,0]^{d}$, with $\ell$ as in (4.13) and where for $k\in R$, the function $\lambda(\,\cdot\,;\Gamma^{(k)}):\mathbb{R}^{\lvert J_{k}\rvert}\to[0,\infty)$ is the density of the exponent measure of a Hüsler–Reiss mev distribution with Gumbel margins.
@@
An expression for $\lambda(\,\cdot\,;\Gamma^{(k)})$ can be obtained for instance from Hentschel et al. (2022, Definition 3.1): writing $d(k):=\max(J_{k})$, we have
@@
and, finally, $\|{\boldsymbol{v}}\|_{M}^{2}=\boldsymbol{v}^{\top}M\boldsymbol{v}$ for $M\in\mathbb{R}^{p\times p}$ and $\boldsymbol{v}\in\mathbb{R}^{p\times 1}$. In fact, replacing $d(k)$ by any other choice of index from $J_{k}$ would give the same function.

### v4

@@ 3 Extreme directions
Upper tail dependence coefficients [7, 37]
@@ 3.1 Definition and characterizations
Recall the exponent measure $\mu$ of $G^{*}$ as defined in (2.4) and $\mathbb{E}=[\boldsymbol{0},\boldsymbol{\infty})\setminus\left\{{\boldsymbol{0}}\right\}$. In this section, let $J\subseteq D$ be a fixed non-empty set and $\mathbb{E}_{J}:=\left\{{\boldsymbol{x}\in\mathbb{E}:\ x_{j}>0\text{ iff }j\in J}\right\}$. The collection $\left\{{\mathbb{E}_{J}:\;\emptyset\neq J\subseteq D}\right\}$ forms a partition of $\mathbb{E}$. The following definition was introduced in Simpson et al. [39] in order to identify extreme directions of the random vector $\boldsymbol{X}$ defined as in (2.3).
@@
$\mathbb{R}_{n,J}^{\epsilon}$ =$=\left\{\boldsymbol{x}\in\mathbb{E}:\ \|\boldsymbol{x}\|>n,\;\boldsymbol{x}_{J}/n\in(\epsilon,\infty)^{J},\;\boldsymbol{x}_{D\setminus J}/n\in[0,\epsilon]^{\lvert D\setminus J\rvert}\right\},$ (3.2)
$\mathbb{C}_{n,J}^{\epsilon}$ =$=\left\{\boldsymbol{x}\in\mathbb{E}:\;\|\boldsymbol{x}\|>n,\;\boldsymbol{x}_{J}/\|\boldsymbol{x}\|\in(\epsilon,\infty)^{J},\;\boldsymbol{x}_{D\setminus J}/\|\boldsymbol{x}\|\in[0,\epsilon]^{\lvert D\setminus J\rvert}\right\},$ (3.3)
@@
The latter sets were introduced in Goix et al. [22, 21] for the $L_{\infty}$-norm. For $\epsilon>0$, the Borel set $\mathbb{R}_{1,J}^{\epsilon}$ is $\mu$-continuous as defined in Section 2. Hence we can make the following link with Proposition 3.2.
@@
$\boldsymbol{M}=\left(\max_{k\in R}\{a_{1k}Z_{k}\},\ldots,\max_{k\in R}\{a_{dk}Z_{k}\}\right).$ (3.9)
@@
$\mathcal{L}(\boldsymbol{U})=\frac{1}{d}\sum_{k\in R}\|\boldsymbol{a}_{\cdot k}\|_{1}\,\delta_{\ln\left({\tfrac{\boldsymbol{a}_{\cdot k}}{\|\boldsymbol{a}_{\cdot k}\|_{1}}}\right)}\left({\cdot}\right),$
with $\delta_{\boldsymbol{x}}(\cdot)$ denoting a unit point mass at $\boldsymbol{x}$. Hence, inequality (6) in Theorem 3.6 is satisfied if and only if $J$ is one of the signatures $J_{k},\,k\in R$.
@@
$\chi_{J}=\mu\left(\left\{\boldsymbol{x}\in\mathbb{E}:\;\boldsymbol{x}_{J}\in(1,\infty)^{J}\right\}\right)=\sum_{\begin{subarray}{c}\bar{J}\subseteq D\\ \bar{J}\supseteq J\end{subarray}}\mu\left(\left\{\boldsymbol{x}\in\mathbb{E}:\;\boldsymbol{x}_{\bar{J}}\in(1,\infty)^{\lvert\bar{J}\rvert},\;\boldsymbol{x}_{D\setminus\bar{J}}\in[0,1]^{\lvert D\setminus\bar{J}\rvert}\right\}\right).$
We have $\chi_{J}>0$ if and only if there exists $\bar{J}\supseteq J$ such that the corresponding term on the right-hand side is positive, and by homogeneity of $\mu$, this is equivalent to $\bar{J}$ being a subset of an extreme direction. We find that $\chi_{J}>0$ if and only if $J$ is a subset of an extreme direction.
@@ 3.2 Density of an mgp distribution with multiple extreme directions
We provide density expressions for potential use in statistical inference on mgp models. As in Section 2.2, let $\boldsymbol{Y}$ be an mgp random vector with distribution $H=\operatorname{GP}(G)$ where $G$ is an mev distribution with Gumbel margins. Further, let $\boldsymbol{T}$ and $\boldsymbol{U}$ be its generators as defined in Section 2. With respect to a dominating measure $\upsilon$, defined as a sum of Lebesgue measures $\lambda_{m}$ of various dimensions $m\in D$, we express the density of $\boldsymbol{Y}$ in terms of the densities of $\boldsymbol{U}$ and $\boldsymbol{T}$. This will be helpful since the latter densities are known for several parametric families, for instance, the logistic and the Hüsler–Reiss model [29, Sections 7.1 and 7.2]. Furthermore, in view of (2.9), the density of the exponent measure $\mu$ on $\left\{{\boldsymbol{x}\geqslant\boldsymbol{0}:\boldsymbol{x}\nleqslant\boldsymbol{1}}\right\}$ is proportional to the mgp density, after appropriate marginal scaling; by homogeneity, this determines the exponent measure density everywhere.
@@
The proofs of Theorems 3.9 and 3.10 are identical to the ones of Theorems 12 and 13 in Rootzén et al. [36]. In the next section, we will propose a model for the $\boldsymbol{U}$-generator that leads to mgp distributions with general extreme directions and whose density can be computed by means of Theorem 3.9.
@@ 4.1 Definition and properties
$\ell_{J_{k}}^{(k)}(\boldsymbol{x}_{J_{k}})=\ell^{(k)}\left({\tilde{\boldsymbol{x}}}\right),\qquad\boldsymbol{x}_{J_{k}}\in[0,\infty)^{J_{k}},$ (4.3)
with $\tilde{\boldsymbol{x}}=\sum_{j\in J_{k}}x_{j}\boldsymbol{e}_{j}$ and $\boldsymbol{e}_{j}$ the $j$-th canonical unit vector in $\mathbb{R}^{d}$. If $J_{k}$ is a singleton, then $\ell_{J_{k}}^{(k)}$ is the identity function on $[0,\infty)$. The following theorem was proved in Stephenson [40, Theorem 1] for the particular case where $J_{1},\ldots,J_{r}$ are all different.
@@
The random vector $\boldsymbol{M}$ in (4.1) follows an mev distribution denoted by $G_{\boldsymbol{M}}$ with unit-Fréchet margins and stdf
@@
In the context of anomaly detection, Chiapino et al. [6] define a mixture model for the angular measure of $G_{\boldsymbol{M}}$ in terms of probability measures on the faces of the unit simplex. In contrast, the mixture model in Definition 4.1 is obtained by an aggregation of lower-dimensional mev distributions.
@@
as defined in Section 2.2. We will refer to $H$ as the mgp distribution associated with the mixture model $\boldsymbol{M}$ and to $\boldsymbol{Y}\sim H$ as the mgp random vector associated with the mixture model $\boldsymbol{M}$. For a non-empty set $J\subseteq D$, recall the set $\mathbb{A}_{J}$, the projection $\pi_{J}$ defined in Eq. (2.1), and the representation $\boldsymbol{U}$ of the mgp distribution $H$ as in Section 2.2. For $k\in R$, since the only extreme direction of the $\mathbb{R}^{J_{k}}$-valued random vector $\boldsymbol{Z}_{J_{k}}^{(k)}$ is $J_{k}$, Theorem 3.6 yields the existence of a generator $\boldsymbol{U}^{(k)}$ taking values in $\mathbb{R}^{J_{k}}$ which satisfies (2.15) and
$\operatorname{\mathsf{E}}\left[{\max_{j\in J_{k}}\left\{{x_{j}\mathrm{e}^{U_{j}^{(k)}}}\right\}}\right]=\ell_{J_{k}}^{(k)}\left({\boldsymbol{x}_{J_{k}}}\right),\qquad\boldsymbol{x}_{J_{k}}\in[0,\infty)^{J_{k}}.$ (4.6)
@@ 4.2.1 Mixture logistic model
$\ell_{J_{k}}^{(k)}(\boldsymbol{y})=\left({\sum_{j\in J_{k}}y_{j}^{1/\alpha_{k}}}\right)^{\alpha_{k}},$ (4.9)
for $\boldsymbol{y}\in[0,\infty)^{J_{k}}$, where $10<\alpha_{k}<$ 1. The only extreme direction of $\boldsymbol{Z}_{J_{k}}^{(k)}$ is $J_{k}$, so that, following Definition 4.1 and Remark 4.2, the mixture logistic model $\boldsymbol{M}$ is well-defined. Let $G_{\boldsymbol{M}}$ denote the cdf of $\boldsymbol{M}$; note that $G_{\boldsymbol{M}}$ is an mev distribution with unit-Fréchet margins. We compute its stdf using Theorem 4.4.
@@
In case the sets $J_{1},\ldots,J_{r}$ are all different, the stdf $\ell$ in (4.10) coincides with the one of the asymmetric logistic model introduced in Tawn [41].
Next we construct a $\boldsymbol{U}$-generator of the mgp distribution $H=\operatorname{GP}(\tilde{G}_{\boldsymbol{M}})$, where $\tilde{G}_{\boldsymbol{M}}$ is the cdf of $\left({\ln M_{1},\ldots,\ln M_{d}}\right)$, an mev distribution with Gumbel margins. Let $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ be random vectors in $\mathbb{R}^{J_{1}},\ldots,\mathbb{R}^{J_{r}}$, respectively, such that
@@ 4.2.2 Mixture Hüsler–Reiss model
For $k\in R$, recall $J_{k}$ in (4.2) and suppose that $\boldsymbol{Z}_{J_{k}}^{(k)}$ has Hüsler–Reiss mev distribution [26] with variogram matrix [^1] A matrix $\Gamma\in\mathbb{R}^{d\times d}$ is a variogram matrix if it is symmetric, has zero diagonal, and satisfies $\boldsymbol{v}^{\top}\Gamma\boldsymbol{v}<0$ for all $\boldsymbol{0}\neq\boldsymbol{v}^{\top}\perp\boldsymbol{1}$. To each positive definite covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$ is associated a variogram matrix $\Gamma\in\mathbb{R}^{d\times d}$ via $\Gamma_{st}=\Sigma_{ss}+\Sigma_{tt}-2\Sigma_{st}$ for $s,t\in D$. $\Gamma^{(k)}\in\mathbb{R}^{J_{k}\times J_{k}}$; see, e.g., Huser and Davison [25]. If $J_{k}$ is a singleton, the stdf $\ell_{J_{k}}^{(k)}$ of $\boldsymbol{Z}_{J_{k}}^{(k)}$ reduces to the identity function on $[0,\infty)$. Otherwise,
$\ell_{J_{k}}^{(k)}(\boldsymbol{y})=\sum_{j\in J_{k}}y_{j}\,\Phi_{\lvert J_{k}\rvert-1}\left({\eta^{j,(k)}\left({\boldsymbol{y}}\right);\Sigma^{j,(k)}}\right),\qquad\boldsymbol{y}\in[0,\infty)^{J_{k}},$
where for $j\in J_{k}$, $\eta^{j,(k)}\left({\boldsymbol{y}}\right)$ is equal to $\left\{{\ln(y_{j}/y_{s})+\Gamma^{(k)}_{js}/2}\right\}_{s\in J_{k};s\neq j}$ with $y_{j}/0=\infty$ if $y_{j}>0$ and $\infty$ if $y_{j}=0$, where the matrix $\Sigma^{j,(k)}$ is defined by
$\Sigma^{j,(k)}:=\frac{1}{2}\left\{{\Gamma_{js}^{(k)}+\Gamma_{jt}^{(k)}-\Gamma_{st}^{(k)}}\right\}_{s,t\in J_{k}\setminus\{j\}}\in\mathbb{R}^{(J_{k}\setminus\{j\})\times(J_{k}\setminus\{j\})}$
@@
To construct a generator of the mgp distribution $H=\operatorname{GP}(\tilde{G}_{\boldsymbol{M}})$ of the mixture Hüsler–Reiss model, let $k\in R$ and let $\tilde{\Sigma}^{(k)}\in\mathbb{R}^{J_{k}\times J_{k}}$ be a positive definite covariance matrix with variogram matrix $\Gamma^{(k)}$; see, e.g., Segers [38] for a possible construction. Let $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ be random vectors in $\mathbb{R}^{J_{1}},\ldots,\mathbb{R}^{J_{r}}$, respectively, such that
@@
for $\boldsymbol{y}\in[-\infty,\infty)^{d}\setminus[-\infty,0]^{d}$, with $\ell$ as in (4.13) and where for $k\in R$, the function $\lambda(\,\cdot\,;\Gamma^{(k)}):\mathbb{R}^{J_{k}}\to[0,\infty)$ is the density of the exponent measure of a Hüsler–Reiss mev distribution with Gumbel margins.
@@
An expression for $\lambda(\,\cdot\,;\Gamma^{(k)})$ can be obtained for instance from Hentschel et al. [23, Definition 3.1]: writing $d(k):=\max(J_{k})$, we have
@@
and, finally, $\|{\boldsymbol{g}}\|_{M}^{2}=\boldsymbol{g}^{\top}M\boldsymbol{g}$ for $M\in\mathbb{R}^{p\times p}$ and $\boldsymbol{g}\in\mathbb{R}^{p\times 1}$. In fact, replacing $d(k)$ by any other choice of index from $J_{k}$ would give the same function.